Circular No. 6855

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J0421+560 D. Smith and R. Remillard, Massachusetts Institute of Technology;

and J. Swank, T. Takeshima, and E. Smith, Goddard Space Flight Center, NASA, report: “The RXTE All-Sky Monitor has detected a bright and rapidly rising x-ray transient, which we designate XTE J0421+560. The best-fit position is α = 4^{h}21^{m}.01, δ = +56^{o}03′.5 (equinox 2000.0), with an error radius (90-percent confidence) of 4′. The ASM countrate (2–12 keV) was below 40 mCrab (at 3σ) on Mar. 31.36 UT, rising rapidly thereafter to 139 (Mar. 31.64), 504 (31.77), and 1880 mCrab (Apr. 1.04). An RXTE PCA observation, which began on Apr. 1.08, confirms the presence of a new source with a flux near 2 Crab. The ASM spectrum appears very hard, but the PCA light curve shows no obvious pulsations at periods longer than 0.25 s. Observations at other wavelengths are urgently needed.”

GRB 980329 A. Guarnieri, Universitá di Bologna; G. Clementini and G. Valentini,

Osservatorio Astronomico di Bologna; A. Castro-Tirado and J. Gorosabel, Laboratorio de Astrofisica Espacial y Fisica Fundamental, Madrid; and A. Pedrosa, Centro de Astrofisica da Universidade do Porto, report on behalf of a large collaboration: “R-band images containing the error box for 1SAX J0702.6+3850 (IAUC 6854) were obtained on Mar. 29.79 and 30.81 UT at the 1.52-m telescope at Bologna Observatory, on Mar. 29.90, 30.91, and 31.95 at the 1.0-m Jacobus Kapteyn Telescope at La Palma, and on Mar. 31.02 and 31.92 at the OAN 1.52-m telescope at the German-Spanish Calar Alto Observatory. No variable object is found down to R = 22; any variation was smaller than 0.3 mag. Observations at other wavelengths, especially infrared, are encouraged.”

NOVA SAGITTARII 1998 Further visual magnitude estimates: Mar. 27.33 UT, 8.4 (J. G. de

S. Aguiar, Campinas, Brazil); 30.138, 9.0 (K. Hornoch, Lelekovice, Czech Republic); 30.177, 9.2 (P. Schmeer, Bischmisheim, Germany); 30.213, 9.2 (A. Pereira, Cabo da Roca, Portugal); 31.124, 8.9 (Hornoch); 31.130, 9.0 (M. Plsek, Lelekovice, Czech Republic); Apr. 1.194, 8.8 (Pereira).

1998 April 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green